OpenStudy (anonymous):

simplify 169/12 Show work.

OpenStudy (anonymous):

14.0833333

OpenStudy (matt101):

To simplify a fraction, divide the numerator and denominator by the same integer. However, in this case, there are no common factors between the two numbers (169 = 13*13, 12 = 2*2*3), so the fraction is already in its simplest form.